OpenStudy (crazycocob):

Please help!!! 1.Simplify. 5x+1=2x+7 2.simplify. 7x+3=5-2x Which equation has infinity many solution? Please show your work. A.3x-5=3(5+x) B.(x+3)=3(x+5) C.3x-5=3(5-x) D.(x+3)=-5(3-x)

OpenStudy (crazycocob):

I give medals!

OpenStudy (anonymous):

1). x = 2

jigglypuff314 (jigglypuff314):

5x + 1 = 2x + 7 subtract 2x from both sides subtract 1 from both sides divide both sides by 3 7x + 3 = 5 - 2x add 2x to both sides subtract 3 from both sides divide both sides by 9 and simplify

jigglypuff314 (jigglypuff314):

the last one seems a bit odd... the way you would solve it would be solving for all options so for example 3x - 5 = 3(5 + x) distribute so 3x - 5 = 15 + 3x then subtract both sides by 3x -3x -3x but then you're left with -5 = 15 but negative 5 does not equal 15 therefore A would be considered "no solution" to get "infinitely many solutions" you would have to end up with something like 5 = 5